Consider a binomial experiment with

n = 7 trials

where the probability of success on a single trial is

p = 0.40.

(Round your answers to three decimal places.)

(a) Find

P(r = 0).




(b) Find

P(r ≥ 1)

by using the complement rule.
Answer #1

(a) P(r=0)

=0.028

(b)

P(r ≥ 1)

=1-P(r=0)

=1-0.028...............................from (a)

=0.972